# Intermittent time series forecasting with Gaussian Processes and Tweedie likelihood

### v4

@@ Appendix D Data and code availability
The data sets used in the experiments are downloadable by running first the file datasets.R in R and then the Python notebook datasets.ipynb. The files are located in the data folder at the GitHub page of the project, which will be released upon acceptance.
The code used for the experiments will be fully available at the GitHub link. The folder “tutorial” contains a simple usage example for our model.
The implementation of our GP models is based on GPyTorch (Gardner et al., 2018). Our contribution to public libraries is given by an implementation of the Tweedie class in the distributions module of PyTorch (Paszke et al., 2019), and the classes TweedieLikelihood and NegativeBinomialLikelihood in the likelihoods module of GPyTorch. Upon acceptance we will contribute to those packages with pull requests.
The GitHub project also contains intermittentGP class. It has methods to build, train and make predictions under the specification of the likelihood, the scaling and different traing hyperparameters, such as the number of epochs or the learning rate of the optimizer.

### v5

@@ Appendix D Data and code availability
The data sets used in the experiments are downloadable by running the files located in the “data” folder at the GitHub page of the project, https://github.com/StefanoDamato/TweedieGP. The code for reproducing the experiments, tables, and plots in this paper is fully available too. In particular, our models are implemented in the intermittentGP class. It has methods to build, train and make predictions under the specification of the likelihood, the scaling and different training hyper-parameters, such as the number of epochs or the learning rate of the optimizer. The “tutorial” folder contains a simple usage example for our models.
The implementation of our GP models is based on GPyTorch (Gardner_Pleiss_Weinberger_Bindel_Wilson_2018). Our contribution to public libraries is given by an implementation of the Tweedie class in the distributions module of PyTorch (pytorch_2019), and the classes TweedieLikelihood and NegativeBinomialLikelihood in the likelihoods module of GPyTorch. We will contribute to those packages with pull requests.